OpenStudy (anonymous):

We are interested in the dimensions of a certain rectangle. This rectangle has length twice the side of the square and width three units less than the side of this square. If the two areas are equal, what are the rectangles dimensions (w x h)?

OpenStudy (anonymous):

let the length of the side of the square be x then the area of the square =x^2 now length of the rectangle = 2x and width of rectangle = x-3 hence area of the rectangle = 2x(x-3) as per the question we have 2x(x-3) = x^2 now solve the resulting quadratic..

OpenStudy (anonymous):

i got x = -4

OpenStudy (anonymous):

2x^2 -6x =x^2 2x^2 -6x -x^2=0 x^2-6x=0 x(x-6) =0 either x= 0 or x=6 we neglect x= 0 (in this case square doesn't exist) so take x=6

OpenStudy (anonymous):

so lenth of the rectangle =2*6 = 12 and width of the rectangle= 6-3= 3

OpenStudy (anonymous):

ohhh! okay thank you for explaining it to me!

OpenStudy (anonymous):

welcome dear